One pound of saturated steam at 689.5 kPa is compressed in a reversible non-flow process from the specific volume of 0.2768 m?/kg to a specific volume of 0.001107 m3/kg. If the pressure remains constant at 689.5 kPa and the internal energy decreases 1878.2 kJ/kg during the process, determine how much heat will be transferred and whether it will be added to our abstracted from the working substance .
1
Expert's answer
2020-03-24T09:57:37-0400

"1" pound "=" "0.45359237" "kg"


"Q=\Delta U+p\Delta V\to \frac{Q}{m}=\frac{\Delta U}{m}+p\frac{\Delta V}{m}"


"q_m=-1878.2+689.5\cdot(0.001107-0.2768)\approx-2069\frac{kJ}{kg}"


"q=-2069\cdot0.45359237\approx-938.5kJ"

The heat will be transferred from the working substance.